# Evaluating multiple models using labeled and unlabeled data

### v2

@@ 7 Discussion
In this paper, we presented Semi-Supervised Model Evaluation (SSME), a method which supplements sparse labeled data with unlabeled data to more accurately estimate classifier performance. SSME exploits three aspects of the current machine learning landscape: (i) there are frequently multiple classifiers for the same task, (ii) continuous classifier scores are often available for all classes, and (iii) unlabeled data is often far more plentiful than labeled data. We show that across multiple tasks, architectures, and modalities SSME substantially outperforms using labeled data alone and standard baselines.
These results suggest several directions for future work. First, each of the metrics we examined center on evaluating a single classifier. But because SSME estimates the full joint distribution $P(y,\mathbf{s})$, it could also be used to to measure properties of the classifiers as a set. For instance, recent work has highlighted the importance of measuring systemic failures (Kleinberg & Raghavan, 2021) where all classifiers produce errors on the same instances. Second, our experiments assess settings in which the unlabeled data is sampled from the same distribution as the labeled data. Although this is common — for example, when a random subset of examples is annotated — there are other settings where the available unlabeled data systematically differs from the labeled data (Sagawa et al., 2021). Applying SSME to those settings represents a natural direction for future work. Finally, future work could also extend SSME beyond classification to estimate the joint distribution $P(y,\mathbf{s})$ for continuous $y$ (e.g, using a mixture density network) or structured $y$ (e.g. graphs) (Vishwakarma & Sala, 2022). More generally, our results strongly indicate that when large amounts of labeled data are unavailable, semi-supervised evaluation can be valuable. SSME is one way to do this, but other approaches are worth exploring.
Acknowledgments
We thank Gabriel Agostini, Serina Chang, Erica Chiang, Sidhika Balachandar, Victor Butoi, Zhi Liu, Katie Matton, Maz Abulnaga, David Li, and Nikhil Garg for their constructive feedback and discussions. This work was supported by an NSF Graduate Research Fellowship under grant number 2141064, a Hertz Fellowship, Quanta Computer, Wistron Corporation, NIH 1R35GM141861, NIH R01 HG010959, Google Research Scholar award, NSF CAREER #2142419, Optum, a CIFAR Azrieli Global scholarship, a gift to the LinkedIn-Cornell Bowers CIS Strategic Partnership, and the Abby Joseph Cohen Faculty Fund. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funders.
@@ Appendix A Related work
Unsupervised performance estimation
involves estimating the performance of a model given only unlabeled data. Methods designed to address this problem often focus on out-of-distribution samples, where labeled data is scarce and model performance is known to degrade. Several works have illustrated strong empirical relationships between out-of-distribution generalization and thresholded classifier confidence (Garg et al., 2022), dataset characteristics (Deng & Zheng, 2021; Guillory et al., 2021), in-distribution classifier accuracy (Miller et al., 2021), and classifier agreement (Parisi et al., 2014; Platanios et al., 2017; Baek et al., 2022).
[…]
Our work is also similar, in spirit, to methods that learn to debias classifier predictions on a small set of labeled data and then apply that debiasing procedure to classifier predictions on unlabeled examples. Prediction-powered inference (Angelopoulos et al., 2023) and double machine learning (Chernozhukov et al., 2018) both learn a debiasing procedure to ensure that unlabeled metric estimates (e.g., accuracy) are statistically unbiased. One of the baselines we compare to, AutoEval (Boyeau et al., 2024), is built atop prediction-powered inference.
@@ B.1 Real datasets and classifier Sets
We provide additional detail for the six datasets we use in our work, including ground truth $p(y)$ for each dataset and ground truth metrics for each classifier in the associated classifier set in Table S1 and Table S2. As discussed, each dataset is split into a training split (provided to each classifier as training data), an estimation split (provided to each performance estimation method), and an evaluation split (used to compute ground truth metrics for each classifier). We determine training splits based on prior work. We then split the remaining data in half (randomly, for each run) to produce the estimation and evaluation splits. We then subsample the estimation split to have $n_{l}$ labeled examples and $n_{u}$ unlabeled examples. We ensure that the labeled data always includes at least one example from each class. Thus, the estimation split contains $n_{l}+n_{u}$ examples in each experiment, and the evaluation split for each task is fixed across runs (exact sample sizes reported below).
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
  6. 6.
[…]
@@
Dataset Classifier Acc ECE
AG News all-MiniLM-L12-v2 84.8 4.2
mxbai-embed-large-v1 85.0 14.4
multi-qa-MiniLM-L6-cos-v1 85.6 5.2
bge-small-en-v1.5 85.2 16.9
bge-large-en-v1.5 86.8 4.8
bge-base-en-v1.5 86.6 5.6
all-mpnet-base-v2 86.7 2.9
all-MiniLM-L6-v2 83.8 3.8
paraphrase-multilingual-MiniLM-L12-v2 85.1 9.6
paraphrase-MiniLM-L6-v2 86.0 8.9
MultiNLI distilbert-SqrtReWeight 81.4 9.2
@@
distilbert-ReSample 81.4 8.2
distilbert-IRM 64.8 6.1
ImagenetBG ResNet-ReWeight 86.6 7.8
ResNet-ReSample 87.4 7.7
ResNet-Mixup 88.6 7.7
ResNet-IRM 54.1 30.9
Table S2: Ground truth classifier metrics on multiclass tasks. We report ground truth performance for classifiers in the sets associated with each multiclass task. Each of the LLMs fine-tuned for AG News are sentence transformers, while the MultiNLI classifiers all use DistilBERT (Sanh, 2019) as the base architecture. The base architecture on ImagenetBG is a ResNet-50.
@@ B.2 Baselines
  • Pseudo-Labeling: We train a logistic regression with the default parameters associated with the scikit-learn implementation (Pedregosa et al., 2011). Experiments with alternative function classes (e.g. a KNN) revealed no significant differences in performance.
@@
  • Bayesian-Calibration: Bayesian-Calibration operates on each classifier individually. We make use of the implementation made available by Ji et al. (2020). Extending the proposed approach to multi-class tasks is not straightforward, so we compare to Bayesian-Calibration only on binary tasks.
@@
  • Active-Testing: We implement Active-Testing, where the method selects a fixed number of examples to label out of a pool of unlabeled examples, according to the approach proposed by Kossen et al. (2022). We select examples according to the acquisition strategy for estimating accuracy, a metric for which a public implementation is available, and limit our comparison to this metric.
  • AutoEval: We implement AutoEval using an implementation made available by the authors (Boyeau et al., 2024). The implementation, to the best of our knowledge, only supports accuracy estimation across a set of classifiers, so we limit our comparison to this metric.
B.3 Semisynthetic dataset and classifier sets
As with the real datasets, we produce three splits: a training split to learn the classifiers (50 examples), an estimation split for the performance estimation methods (20 labeled examples and 1000 unlabeled examples), and an evaluation split to measure ground truth values for each metric (10,000 examples). Each classifier is a logistic regression with default L2 regularization.
B.4 Computing effective sample size
Appendix C Normalizing Flow
@@
$\mathbf{Z}|(Y=k)\sim\mathcal{N}(\mu_{k},\Sigma_{k})$
Thus, the marginal distribution of $\mathbf{Z}$ is $p_{\mathbf{Z}}(\mathbf{z})=\sum_{k=1}^{K}\mathcal{N}(\mathbf{z}|\mu_{k},\Sigma_{k})\cdot p(y=k)$ is the overall density of $\mathbf{z}$. We apply our invertible transformation $f_{\theta}$ to obtain $\mathbf{s}=f_{\theta}(\mathbf{z})$. To find $p(\mathbf{s}|y=k)$, we follow the approach of Izmailov et al. (2020):
We use the RealNVP architecture (Dinh et al., 2016) to parameterize $f_{\theta}$ using 10 coupling layers, 3 fully-connected layers, and a hidden dimension of 128 between the fully connected layers. Our normalizing flow is lightweight and trains in less than a minute for each dataset in our experiments section using 1 80GB NVIDIA A100 GPU.
@@ Appendix D Supplementary results
Figure S1: Impact of average accuracy across classifiers in set on SSME’s performance.
Figure S2: Impact of average ECE across classifiers in set on SSME’s performance.
D.1 Results reporting mean absolute error

### v3

@@ 7 Discussion
In this paper, we presented Semi-Supervised Model Evaluation (SSME), a method that supplements limited labeled data with unlabeled data to more accurately estimate classifier performance. SSME leverages three features of the current machine learning landscape: (i) there are frequently multiple classifiers for the same task, (ii) continuous classifier scores are often available for all classes, and (iii) unlabeled data is often far more plentiful than labeled data. We provide theoretical bounds on performance estimation error that help explain SSME’s advantages over prior work. We show that across multiple tasks, architectures, and modalities, SSME substantially outperforms using labeled data alone and standard baselines.
Our results, and their accompanying limitations, suggest several directions for future work. First, each of the metrics we examined (e.g., AUC) evaluate a single classifier. But because SSME estimates the full joint distribution $P(y,\mathbf{s})$, it could be used to measure properties of the classifiers as a set. For instance, recent work has highlighted the importance of measuring algorithmic monoculture (kleinberg_algorithmic_2021, ) where all classifiers produce errors on the same instances. Second, our experiments assess settings in which the unlabeled data is sampled from the same distribution as the labeled data. Although this is common — for example, when a random subset of examples is annotated — many real-world scenarios violate this assumption, and extending SSME to settings with such distribution shift represents an important next step (sagawa2021extending, ). Finally, our current implementation relies on kernel density estimation, which can scale poorly in high-dimensional spaces (e.g., with many classifiers or classes). Exploring more scalable density estimators could further broaden SSME’s applicability. More generally, our results indicate that SSME offers a principled and practical framework for accurate performance estimation in the absence of a large labeled dataset.
@@ Appendix A Related work
Unsupervised performance estimation involves estimating the performance of a model given only unlabeled data. Methods designed to address this problem often focus on out-of-distribution samples, where labeled data is scarce and model performance is known to degrade. Several works have illustrated strong empirical relationships between out-of-distribution generalization and thresholded classifier confidence [28], dataset characteristics [21, 30], in-distribution classifier accuracy [48], and classifier agreement [52, 57, 5].
Several works have formalized when unsupervised model evaluation is possible [23, 13, 28, 47, 73, 20, 62], and propose assumptions under which estimates of performance are recoverable. [23] and [6] assume knowledge of $p(y)$ in the unlabeled sample. [64] assume conditionally-independent subsets of the observed features, inspired by conditional-independence assumptions made in works such as [18]. [30] assume classifier calibration on unlabeled samples. [13] assume a sparse covariate shift model, in which a subset of the features’ class-conditional distribution remains constant. [47] illustrate misestimation of $p(y)$ in the unlabeled example, and assume that $p(y)$ out-of-distribution is close to $p(y)$ in-distribution. As [28] highlight, assumptions are necessary to make any claim about the nature of unsupervised model evaluation, and the above methods are a representative sample of assumptions made Finally, there has been a surge of interest in unsupervised performance estimation in the context of large language models [77, 34]. A standard approach here is to use a large language model to adjudicate the quality of text generated by other language models. Methods in this literature are often specific to large language models, while SSME is not.
Our work is also similar, in spirit, to methods that learn to debias classifier predictions on a small set of labeled data and then apply that debiasing procedure to classifier predictions on unlabeled examples. Prediction-powered inference [3] and double machine learning [14] both learn a debiasing procedure to ensure that unlabeled metric estimates (e.g., accuracy) are statistically unbiased. One of the baselines we compare to, AutoEval [11], is built atop prediction-powered inference.
@@ B.1 Real datasets and classifier sets
We provide additional detail for the five datasets we use in our work, including ground truth $p(y)$ for each dataset and ground truth metrics for each classifier in the associated classifier set in Table S1 and Table S2. As discussed, each dataset is split into a training split (provided to each classifier as training data), an estimation split (provided to each performance estimation method), and an evaluation split (used to compute ground truth metrics for each classifier). We determine training splits based on prior work. We then split the remaining data in half (randomly, for each run) to produce the estimation and evaluation splits. We then subsample the estimation split to have $n_{l}$ labeled examples and $n_{u}$ unlabeled examples. We ensure that the labeled data always includes at least one example from each class. Thus, the estimation split contains $n_{l}+n_{u}$ examples in each experiment, and the evaluation split for each task is fixed across runs (exact sample sizes reported below). No performance estimation method sees data from the evaluation split, which is used to evaluate the performance estimates.
@@
    MIMIC-IV: We use three binary classification tasks from MIMIC-IV [38], a large dataset of electronic health records describing 418K patient visits to an emergency department. We focus on three tasks: hospitalization (predicting hospital admission based on features available during triage, $p(y=1)=0.45$), critical outcomes (predicting inpatient mortality or a transfer to the ICU within 12 hours, $p(y=1)=0.06$), and emergency department revisits (predicting a patient’s return to the emergency department within 3 days, $p(y=1)=0.03$). We split and preprocess data according to prior work [72, 50]. No patient appears in more than one split. For each task, the evaluation split contains 70,439 examples. The classifiers in the associated set differ by function class (logistic regression, decision tree, and multi-layer perceptron) and random seed (0, 1, 2).
@@
    Toxicity detection: The task is to predict presence of toxicity given an online comment, using data from CivilComments [9, 42] where $p(y=1)=0.11$. The evaluation split contains 66,891 examples. The classifiers in the associated set differ by training loss (ERM, IRM, and CORAL) and random seed (0, 1, 2).
@@
    Biochemical property prediction The task is to predict presence of a biochemical property based on a molecular graph, using data from the Open Graph Benchmark [33]. We focus on the task of predicting whether a molecule inhibits SARS-CoV virus maturation, where $p(y=1)=0.09$. We filter out examples for which no label is observed (i.e. the molecule was not screened at all) because it is impossible to evaluate our performance estimates on those examples. Doing so reduces data held-out from training from 43,793 to 28,325 examples. The evaluation split then contains half, or 14,163, of those examples. The classifiers in the associated set differ by training loss (ERM, IRM, and CORAL) and random seed (0, 1, 2).
@@
    News classification The task is to predict one of four news types based on the title and description of an article [75]. The dataset contains 7,600 examples, with 2,550 examples reserved for the evaluation split. The classes are balanced. Classifiers differ by the base LLM used to perform news classification. The first LLM is Llama-3.2-3B-Instruct, and we obtain probabilities over all classes through zero-shot prompting (i.e. proving one example at a time, accompanied by options) using code provided by [74]. The three remaining LLMs are closed-source models provided by OpenAI: text-embedding-3-small, text-embedding-3-large, and text-embedding-ada-002. We train a classifier atop these embeddings by training a multiclass logistic regression on 2,500 embeddings of labeled examples.
@@
    Sentence classification The task is to predict one of three textual entailments from a sentence [70]. The classes are balanced and the evaluation split contains 61,856 examples. Classifiers differ by training loss (ReWeight, ReSample, IRM, and SqrtReWeight) according to [73].
@@
dataset model Acc ECE
AG News llama-3.2-3B-Instruct 85.6 8.6
text-embedding-3-large 90.1 8.8
text-embedding-3-small 89.8 9.1
text-embedding-ada-002 89.0 9.3
MultiNLI distilbert-IRM 64.8 6.1
distilbert-ReSample 81.4 8.2
@@
distilbert-SqrtReWeight 81.4 9.2
Table S2: Ground truth classifier metrics on multiclass tasks. We report ground truth performance for classifiers in the sets associated with each multiclass task. Each of the LLMs fine-tuned for AG News are sentence transformers, while the MultiNLI classifiers all use DistilBERT [63] as the base architecture.
@@ B.2 Baselines
  • Pseudo-Labeling (PL): We train a logistic regression with the default parameters associated with the scikit-learn implementation [54]. Experiments with alternative function classes (e.g. a KNN) revealed no significant differences in performance.
@@
  • Bayesian-Calibration (BC): Bayesian-Calibration operates on each classifier individually. We make use of the implementation made available by [37]. Extending the proposed approach to multi-class tasks is not straightforward, so we compare to Bayesian-Calibration only on binary tasks. We compare to Bayesian-Calibration on binary tasks, as it does not extend naturally to multiclass settings.
  • SPE: SPE [69] ise a semi-supervised single classifier evaluation method that relies on parametric assumptions about the distribution of classifier scores. They find that the truncated Normal distribution demonstrates reasonable fit across datasets; accordingly, we implement SPE as a mixture of truncated Normal distributions fit to each classifier’s predictions individually.
  • Dawid-Skene: We implement Dawid-Skene with a tolerance of 1e-5 and a maximum number of EM iterations of 100 (the default parameters), using the following public implementation: https://github.com/dallascard/dawid_skene. Dawid-Skene accepts discrete predictions, so we discretize classifier predictions using thresholding the predicted class probability at $\frac{1}{K}$.
  • Active-Testing: We implement Active-Testing, where the method selects a fixed number of examples to label out of a pool of unlabeled examples, according to the approach proposed by [43]. We select examples according to the acquisition strategy for estimating accuracy, a metric for which a public implementation is available, and limit our comparison to this metric.
  • AutoEval: We implement AutoEval using an implementation made available by the authors [11]. The implementation, to the best of our knowledge, only supports accuracy estimation across a set of classifiers, so we limit our comparison to this metric. We compare to AutoEval on accuracy estimation, as additional metrics are not supported by the public implementation.
B.3 Computing effective sample size
Appendix C Proof of Theorem 1
We derive a high-probability error bound on performance estimates based on prior results in semi-supervised mixture models [67]. The proof consists of two steps: we first bound the error in estimation of component separation. We then use the relationship between component separation and classifier performance to bound the error in estimating classifier performance.
Model of $y$ and $\mathbf{s}$.
We consider a stylized binary classification setting drawn from previous work [67] where classifier scores $\mathbf{s}$ are drawn from a balanced mixture of two Gaussians. For data in the positive class, classifier scores are drawn from a multivariate Gaussian The Gaussian assumption holds when the distribution of classifier logits follows a normal distribution. $\mathcal{N}(\mathbf{c},I)$, where $\mathbf{c}\in\mathbb{R}^{d}$. For data in the negative class, classifier scores are drawn from a multivariate Gaussian $\mathcal{N}(0,I)$. Let $n_{u}$ refer to the number of unlabeled examples, $n_{\ell}$ the number of labeled examples, and $d$ the number of classifiers. Additional classifiers increase the dimensionality of the multivariate Gaussian.
Estimator
We reason about SSME’s performance using a simplified semi-supervised learning algorithm. The estimator, UL+, first uses the unlabeled data to identify decision boundaries, and then uses the labeled data to assign decision boundaries to classes. As the authors note, this is not the most efficient use of labeled data. We expect that using the labeled data in the expectation-maximization procedure – as SSME does – to weakly outperform UL+ because it uses labeled data when learning the decision boundaries. For additional detail on the estimator, please refer to Section 2.3 of [67].
Theorem 1
Let $\widehat{AUC}_{k}$, $\widehat{ACC}_{k}$ be estimators for $AUC_{k}$, $ACC_{k}$, estimated using UL+, and $\mathbf{c}_{k}$ be the L2 distance between components along dimension $k$. Assume $n_{u}\gtrsim\max\left\{d,\frac{d}{\mathbf{c}^{4}},\frac{\log(n_{u})}{\min\{\mathbf{c}^{2},\mathbf{c}^{4}\}},\frac{d\log(dn_{u})}{\mathbf{c}^{6}}\right\}$ and $d\geq 2$. With probability 1p1-p, errors in the estimated AUC and accuracy of classifier $k$ differ by at most:
$|AUC_{k}-\widehat{AUC}_{k}|\leq$ $\quad\Phi\big(\frac{\mathbf{c}_{k}}{\sqrt{2}}\big)-\Phi\big(\frac{\mathbf{c}_{k}-\epsilon_{\mathbf{c}}}{\sqrt{2}}\big)$ (5)
$|ACC_{k}-\widehat{ACC}_{k}|\leq$ $\quad\Phi\big(\frac{\mathbf{c}_{k}}{2}\big)-\Phi\big(\frac{\mathbf{c}_{k}-\epsilon_{\mathbf{c}}}{2}\big)$ (6)
[…]
1. Bound error in separation estimation as a function of $n_{l}$, $n_{u}$, $d$, and separation $\mathbf{c}$.
Prior work [67] has established bounds on the error in mean estimation $\epsilon_{\mu}$ in the context of a semi-supervised learning algorithm which first uses the unlabeled data to generate decision boundaries, and uses the labeled data to assign labels to regions. The authors prove mean estimation error can be bounded by:
[…]
where $C_{0}$ is some universal constant and not equivalent to $\mathbf{c}||$. We can convert this expectation bound into a high-probability bound using Markov’s inequality:
[…]
Because we pin $\mu_{0}=0$, error in mean estimation is equal to error in separation estimation in this setting, i.e. $\epsilon_{\mu}=\epsilon_{\mathbf{c}}$. So we can state that, with probability 1p1-p, the mean estimation error obeys the following inequality:
[…]
2. Bound error in performance estimation.
The AUC of the optimal linear classifier (separating classifier predictions for the negative class from classifier predictions in the positive class) is a function of the Mahalanobis distance $\mathbf{c}$, equivalent to the L2 distance in this setting, as prior work has shown [55]:
[…]
where $\Phi$ is the cumulative distribution function of the standard normal distribution. Similarly, the accuracy of the optimal linear classifier is:
[…]
Above, we derived a bound on the error of separation estimation, $\epsilon_{c}$. Now, we map the error in separation estimation to errors in performance estimation (AUC and accuracy). We can state that the following inequality holds:
[…]
Thus:
@@ Appendix C Proof of Theorem 1
Given that we assume the classifiers are better than random (i.e. $\mathbf{c}||$ > 0) the term on the far left of the inequality is larger in magnitude than the term on the far right, so we can simplify as:
[…]
Finally, error in estimated separation is necessarily larger than the error in any one dimension, allowing us to state the inequality for each dimension of the mixture. We can also follow the same logic above to derive a bound on error in estimated accuracy. We can state:
$|\widehat{AUC}_{k}-AUC_{k}|\leq$ $\quad|\Phi\big(\frac{\mathbf{c}_{k}}{\sqrt{2}}\big)-\Phi\big(\frac{\mathbf{c}_{k}-\epsilon_{\mathbf{c}}}{\sqrt{2}}\big)|$ (12)
$|\widehat{ACC}_{k}-ACC_{k}|\leq$ $\quad|\Phi\big(\frac{\mathbf{c}_{k}}{2}\big)-\Phi\big(\frac{\mathbf{c}_{k}-\epsilon_{\mathbf{c}}}{2}\big)|$ (13)
where the following holds with probability 1p1-p:
[…]
We plot bounds in estimated classifier performance as a function of $\epsilon_{\mathbf{c}}$ in Fig. S1.
Figure S1: Estimation error based on original classifier set performance. We plot estimated classifier performance as a function of original classifier performance for various values of $\epsilon_{\mathbf{c}}$, for accuracy (left) and AUC (right). The less accurate the classifier set, the larger the error bands.
C.1 Impact of additional classifiers
[…]
The norm of the augmented vector $\mathbf{c}^{\prime}=[\mathbf{c};\delta]$ exceeds $\sqrt{d+1}$ if and only if
$\delta>\sqrt{d+1-\|\mathbf{c}\|^{2}}.$
If $\|\mathbf{c}\|>\sqrt{d+1}$, then any $\delta>0$ will result in $\|\mathbf{c}^{\prime}\|>\sqrt{d+1}$. Otherwise, $\delta$ must be large enough to compensate for the square root of the difference between the new dimensionality ($d+1$) and the squared norm ($\mathbf{c}||^{2}$). At an extreme, if the new classifier is perfectly correlated with another, it increases dimensionality at no benefit to separation between components, resulting in a looser bound.
@@ Appendix D Supplementary results
D.1 Results on synthetic data
We corroborate theoretical findings from Section 3.2 in synthetic experiments, including settings where the assumptions do not hold. We draw classifier scores for positive examples from distributions centered on a vector $\mathbf{c}$, where entries of $\mathbf{c}$ are drawn from $N(c,.2)$, enforcing a minimum entry of 0.01 (to ensure that each classifier exhibits performance better than random). We consider $n_{u}\in{50,100,250,500,1000,2000}$, $d\in{2,4,6,8,10}$, and $c||\in{0.75,1.0,1.25,1.5}$, and sample 50 datasets for each setting. We provide 20 labeled examples to SSME across all synthetic experiments.
The three theoretical findings we make (using a simplified semi-supervised learning algorithm and a Gaussian assumption) hold in our empirical setting. Our synthetic setting differs from our theoretical setting in two key respects: we use SSME to infer performance and we simulate violations of the Gaussian assumption. Our theoretical findings hold in this more realistic setting; Figure S2 plots results.
D.2 Results reporting mean absolute error